Investigating - Some customers who are being migrated over from Moonfruit may currently be unable to access their emails.
We recommend that affected customers follow the step listed below in order to alleviate this problem.

Log into Webmail (https://webmail.moonfruit.domains) using your current Moonfruit email username and password. You will be able to automatically import all of your existing emails into your new names.co.uk Webmail with the click of a button once you are logged in.

The guide linked below will walk you through this process:

https://www.moonfruit.domains/knowledge/2847.html

If you are still experiencing problems with email retrieval once you have done this, then we recommend using the 'Technical Support' section of your names.co.uk Online Control Panel to send in a support ticket, and we will investigate further.
May 10, 08:27 BST

About This Site

Service status information and Network announcements.

Managed Servers ? Operational
DNS ? Operational
Email - POP3/IMAP Access ? Operational
Linux Hosting ? Operational
Windows Hosting ? Operational
Network ? Operational
WebMail ? Operational
Control Panel ? Operational
Managed Virtual Servers ? Operational
Phone System ? Operational
FTP (Linux Hosting) ? Operational
FTP (Windows Hosting) ? Operational
Email - SMTP Authentication ? Operational
Operational
Degraded Performance
Partial Outage
Major Outage
Maintenance
Past Incidents
Jun 23, 2017

No incidents reported today.

Jun 22, 2017

No incidents reported.

Jun 21, 2017
Completed - This maintenance has been cancelled and will be rescheduled.
Jun 21, 16:15 BST
Scheduled - On June 21st 2017 starting at 2100 BST through to June 22nd 2017 at 0600 BST maintenance will take place on our network between our Reading datacentre and our Telehouse PoP. During this time we will re-route traffic via another path within our network in order to minimise disruption to customers. There is ample capacity elsewhere in our network to support this traffic. No impact is expected but this is an 'at risk' period for our network.
Jun 13, 15:27 BST
Completed - This maintenance period is now being marked as complete, there will be no further works today on the STS.
Jun 21, 14:39 BST
Verifying - The maintenance has been completed and we have moved off of the maintenance bypass.

We will monitor the situation for another hour before marking as complete.
Jun 21, 12:51 BST
Update - This maintenance is now underway and we the STS has transitioned into bypass mode.
Jun 21, 12:29 BST
In progress - Scheduled maintenance is currently in progress. We will provide updates as necessary.
Jun 21, 09:00 BST
Scheduled - On 21st June from 09:00 to 17:00 BST engineers are performing maintenance to an STS power switch system that supplies a proportion of our Reading data centre.

During the window it will be put into maintenance bypass. Services will still be protected by the A side UPS system but unable to transfer to the B side UPS system.

No interruption is anticipated but it is an ‘at risk’ period with reduced level of redundancy.
Jun 7, 10:04 BST
Jun 20, 2017

No incidents reported.

Jun 19, 2017

No incidents reported.

Jun 18, 2017

No incidents reported.

Jun 17, 2017

No incidents reported.

Jun 16, 2017

No incidents reported.

Jun 15, 2017

No incidents reported.

Jun 14, 2017

No incidents reported.

Jun 13, 2017

No incidents reported.

Jun 12, 2017

No incidents reported.

Jun 11, 2017

No incidents reported.

Jun 10, 2017

No incidents reported.

Jun 9, 2017

No incidents reported.